Paynesville Corporation manufactures and sells a preservative used in food and drug manufacturing. The company carries no inventories. The master budget calls for the company to manufacture and sell 110,000 liters at a budgeted price of $150 per liter this year. The standard direct cost sheet for one liter of the preservative follows.

Direct materials (2 pounds @ $9) $ 18
Direct labor (0.5 hours @ $34) 17


Variable overhead is applied based on direct labor hours. The variable overhead rate is $70 per direct-labor hour. The fixed overhead rate (at the master budget level of activity) is $35 per unit. All non-manufacturing costs are fixed and are budgeted at $1.7 million for the coming year.

At the end of the year, the costs analyst reported that the sales activity variance for the year was $480,000 unfavorable.

The following is the actual income statement (in thousands of dollars) for the year.

Sales revenue $ 15,838
Less variable costs
Direct materials 1,738
Direct labor 1,760
Variable overhead 3,480
Total variable costs $ 6,978
Contribution margin $ 8,860
Less fixed costs
Fixed manufacturing overhead 1,100
Non-manufacturing costs 1,280
Total fixed costs $ 2,380
Operating profit $ 6,480


During the year, the company purchased 186,000 pounds of material and employed 45,400 hours of direct labor.


Required:

a. Compute the direct material price and efficiency variances.
b. Compute the direct labor price and efficiency variances.
c. Compute the variable overhead price and efficiency variances.

(For all requirements, enter your answers in whole dollars. Indicate the effect of each variance by selecting "F" for favorable, or "U" for unfavorable. If there is no effect, do not select either option.)

rev: 11_20_2019_QC_CS-191240
Answer #1
Paynesville Corporation Manufectures
1 First of all we have to calculate actual quantity sold by Enterprise
2 we have not been provided directly with this information but we have been provided Sale Activity Variance
3 Sale activity variance:- The variance is calculated by taking the difference between the actual sales volume at the budgeted mix and the budgeted sales volume and multiplying this by the budgeted price to give a monetary amount.
In simple we can say
Sale activity Variance =(Actuale sale at budget mix - Standard Sale)*Standard price
4 Given Information is as followes
Sale Activity variance:- (-) 480000
Actuale sale = x
Standar Price=150
Now:
(-) 480000 = (x - 110000)*150
(-) 3200 = x - 110000
x = 106800
5 Actuale sale = x = 106800
6

Now we solve require part but before that we will one table to get all data

Amount in $
Particular Standard Standard Actual
1 unit 106800 Units 106800 units
Quantity/Hour Rate Amount Quantity/Hour Rate Amount Quantity/Hour Rate Amount
Material (POUND) 2 9 18 213600 9 1922400 186000 9.3441 1738000
Labour (HOUR) 0.5 34 17 53400 34 1815600 45400 38.767 1760000
Variable Overhead (HOUR) 0.5 70 35 53400 70 3738000 45400 76.652 3480000
7 Required Part
a a. Material Price and efficiency Variance
Material Price variance
(Standard Price - Actual Price)*Actual Quantity
(9 - 9.34)*186000 = $ 63240 (u)
Material Efficiency Variance
(Standard Quantity - Actual Quantity)*Standard Price
(213600 - 186000)*9 = $248400(F)
b Dirct labour Price and Efficiency Variance
Direct Labour Price Variance
(Standard Rate - Actual Rate)*Actual Hours
(34-38.8)*45400 =$ 217920 (U)
Labour Efficeincy Variance
(Standard Hours - Actual Hours)*Standard Rate
(53400 - 45400)*34 = $ 272000 (F)
c Variable Overhead and efficiency Variance
Variable Overhead Price Variance
(Standard Rate - Actual Rate)*Actual Hours
(70 - 76.7)*45400 = $ 304180
Varialbe Overhead Efficiency Variance
(Standard Hours - Actual Hours)*Standard Rate
(53400 - 45400) * 70 = $ 560000
All The Above answer are round off to 2 Decimal so there might be difference due to roundoff